As the number of units manufactured increased from 100 to 200, manufacturing cost (total)
increased from $350 to $650. Assume the given data establish the relationship between cost (y)
and number of units made (x), and assume the relationship is linear. Find the equation of the
relationship

Answer by greenestamps(13214)



A linear equation is of the form

y = mx+b

where m is the slope (rate of change of y with respect to x) and b is the value when x is 0.

With the given information, the cost increased by $650-$350=$300 when the number of units produced increased by 200-100=100. So the rate of change of cost (y) with respect to x (number of units) is 300/100=3.

So the equation is of the form

y = 3x+b

Use one of the two given data points to determine b.

When production was 100, the total cost was $350:

350 = 3(100)+b
350 = 300+b
b = 50

ANSWER: y=3x+50 is the linear equation relating the total cost to the number of units produced.